# Sphere Shaping with Dispersion-Aware Sequence Selection for Long-Haul Transmission Systems

### 21 Dec 2023 v1

@@ V-D WDM Transmission Performance
Finally, we maintain the same system configuration as before but extend it to a $5$-channel WDM transmission. The combined net transmission rate for all five channels is $4.4$ Tbits/s. In our earlier investigation, we observed poor performance for D-E-ESS with a block length of $60$. Consequently, we assess the performance starting with a block length of $108$. In Fig. 13, we see that, as the block length increases, the $\mathrm{SNR_{elec}}$ for ESS gradually decreases. However, the increase in GMI due to limited rate loss compensates for the reduction in nonlinear gain, yielding an optimal block length of $300$ for ESS. Conversely, D-E-ESS better maintains nonlinear gains as the block length increases, and its rate loss continues to decrease. This results in D-E-ESS achieving the maximum GMI at a block length of $300$. Particularly, with $16$ selections, D-E-ESS with a block length of $300$ outperforms ESS by $0.02$ bits/4D-symbol, and it achieves a $0.06$ bits/4D-symbol improvement compared to ESS with a block length of $60$. Lastly, it is important to note that $256$ selections, which require more flipping bits, lead to higher rate loss. The additional nonlinear gain does not fully compensate for the performance degradation caused by the increased rate loss, resulting in a lower GMI than with $16$ selections. In a WDM configuration, the gains are lower than the ones observed in a single-channel system as in WDM, a high number of Cross-Phase Modulation components arise and our selection metric does not account for the energy dispersion of the neighboring channels.
VI Discussion and Conclusion
@@
We have proposed two reduced-complexity shaping schemes based on sequence selection and ESS. The first is based on a sign-independent metric (EDI), and the second is based on a sign-dependent metric (D-EDI). These new schemes achieved higher gains than state-of-the-art ESS schemes over short transmission distances. Noticably, D-E-ESS showed gains even for longer transmission distances and higher baud rates over single-channel and WDM systems even when CPR is used to reduce the impact of non-linear phase rotations. Concurrently, we showed that D-EDI can serve as a prediction metric of the generation of NLI because it showed a negative correlation with the measured performance of long-distance transmission systems. Moreover, D-E-ESS retained a robust performance at moderate complexity (lower number of digital dispersion operations to compute the metric) and outperformed conventional ESS across different block lengths. These results pave the way for further optimizations of shaping schemes for long-distance transmissions and confirm the necessity of using sign-dependent and channel-aware metrics for effectively mitigating nonlinear distortions using sequence selection. Nevertheless, these findings tell us that there is still room for improvement. The current probability shaping approach, based on sequence selection, revolves around three primary considerations:
  1. 1.
    Low-complexity candidate sequence generator: the method aims to maintain a low level of complexity for the generation of candidate sequences, allowing for a wider range of performance fluctuations. This enables a higher probability of selecting the optimal sequence and helps to improve the upper limit of sequence-selection performance.
  2. 2.
    Accurate and low-complexity measurement metric: precise measurements are employed to assess the nonlinear performance of each candidate sequence. These measurements play a crucial role in determining which sequence is selected for transmission.
  3. 3.
    Matching of fixed sign bits and FEC encoding process in PCS schemes: another key consideration is to minimize complexity and reduce latency when matching fixed signs to FEC codewords. This optimization helps in improving the overall efficiency and applicability of the scheme.
In this paper, we improved the first two considerations. First, we used flipping bits to generate candidate sequences, instead of using LUTs and XOR-based bit scrambling methods as in [27]. This approach reduces the complexity of the scheme and avoids the additional memory requirements of LUTs. However, our D-EDI-based sequence selection exhibited almost the same performance improvement as the single-channel SSFM-based selection. This implies that we have reached the single-channel performance limit of sequence selection by flipping bits to generate sequences. To achieve further gains, we need to find better methods for the generation of sequences that exhibit larger possible performance variations, thus improving the upper limit of sequence-selection performance.
@@
Second, we proposed a novel metric D-EDI which shows a negative correlation with the nonlinear performance of the candidate sequence, and helps in achieving the same performance as the SSFM-based selection [27], however at a much lower complexity. Yet, for DMB-based transmissions in which the transmission rate per subcarrier is lower than single-carrier transmissions, the impact of channel memory is limited to a few symbols per span or even less than 1 symbol per span, hence only a few dozen symbols for all spans. Thus, the impact of channel memory can be accounted for by concatenating several DMs to minimize the mutual influence of adjacent selected sequences. As the transmission rates continue to increase, for both multi-carrier or single-carrier transmissions, the channel memory can spread over hundreds or thousands of symbols. At the same time, increasing the length of candidate sequences also increases the complexity of the scheme. Hence, modifications and optimizations of D-EDI to achieve excellent performance predictions under high-rate long-distance transmission while maintaining a reduced complexity is still an open question. Other improvements of the metric may address different dispersion management schemes or take into account inter-channel non-linear effects. In this work, we only considered long-haul dispersion-unmanaged systems where the NLI contributions from different spans can be considered independent and D-EDI gives a good performance prediction by summing up EDIs of the sequence at the start of each span.
Third, in this work, we used the multi-block FEC-independent structure proposed by [27], which fixes the signs of symbols and combines the fixed-signs sequence selection with FEC. However, the delay in this structure tends to be significant due to the encoding process involving multiple blocks, leading in this case to severe latency as the receiver should wait for the last block to be received before decoding. A lower-latency and lower-complexity matching method between fixed signs and FEC codewords is still to be discovered.

### 05 Mar 2024 v2

@@ VI-D WDM Transmission Performance
Finally, we maintain the same system configuration as before but extend it to a $5$-channel WDM transmission. The combined net transmission rate for all five channels is $4.4$ Tbits/s. In our earlier investigation, we observed poor performance for D-SS with a block length of $60$. Consequently, we assess the performance starting with a block length of $108$. In Fig. 12, we see that the nonlinear gain associated with short block lengths in ESS without SS diminishes because the CPR compensates for the increased inter-channel nonlinearity. There is an observed increase in GMI with increasing block length due to the reduced rate loss, yielding an optimal block length of $300$ for ESS. Conversely, D-SS better maintains nonlinear gains as the block length increases and its rate loss continues to decrease. This results in D-SS achieving the maximum GMI at a block length of $300$. Particularly, with $16$ selections, D-SS with a block length of $300$ outperforms ESS by $0.02$ bits/4D-symbol, and it achieves a $0.06$ bits/4D-symbol improvement compared to ESS with a block length of $60$. Lastly, it is important to note that $256$ selections, which require more flipping bits, lead to higher rate loss. The additional nonlinear gain does not fully compensate for the performance degradation caused by the increased rate loss, resulting in a lower GMI than with $16$ selections. In a WDM configuration, the gains are lower than the ones observed in a single-channel system as a high number of cross-phase modulation components arise and our selection metric does not account for the energy dispersion of the neighboring channels. Finally, since these sequences are first selected for a single digital band and then used for all bands, the sequences selected by SSFM-SS are the same as D-SS and exhibit the same performance.
VII Discussion and Conclusion
@@
We have proposed a novel sign-dependent metric: D-EDI, which can serve as a prediction metric for the generation of NLI because it showed a negative correlation with the measured performance of long-distance transmission systems. By applying D-EDI (resp. EDI) in sequence selection, we defined the D-SS (resp. E-SS) scheme. These two schemes achieved higher gains with ESS used for distribution matching than state-of-the-art ESS schemes over short transmission distances. Noticeably, D-SS showed gains even for longer transmission distances and higher baud rates over single-channel and WDM systems even when CPR is used to reduce the impact of nonlinear phase rotations, achieving the same performance as sequence selection based on a full SSFM simulation [27]. Moreover, D-SS retained a robust performance at moderate complexity (lower number of digital dispersion operations to compute the metric) and outperformed conventional ESS across different block lengths. These results pave the way for further optimizations of shaping schemes for long-distance transmissions and confirm the necessity of using sign-dependent and channel-aware metrics for effectively evaluating nonlinear distortions.
@@
Nevertheless, these findings tell us that there is still room for improvement. For DMB-based transmissions in which the transmission rate per subcarrier is lower than single-carrier transmissions, the impact of channel memory is limited to a few symbols per span or even less than 1 symbol per span, hence only a few dozen symbols for all spans. Thus, the impact of channel memory can be accounted for by concatenating several DMs to minimize the mutual influence of adjacent selected sequences. As the transmission rates continue to increase, for both multi-carrier and single-carrier transmissions, the channel memory can spread over hundreds or thousands of symbols. At the same time, increasing the length of candidate sequences also increases the complexity of the scheme. Hence, modifications and optimizations of D-EDI to achieve excellent performance predictions under high-rate long-distance transmission while maintaining a reduced complexity is still an open question. Other improvements of the metric may address different dispersion management schemes or take into account inter-channel nonlinear effects. In this work, we only considered long-haul dispersion-unmanaged systems where the NLI contributions from different spans can be considered independent and D-EDI gives a good performance prediction by summing up EDIs of the sequence at the start of each span.